# Information-Theoretic Thresholds for Bipartite Latent-Space Graphs Under Noisy Observations

### 11 Mar 2026 v3

@@ 5. Information-Theoretic Hardness
$\frac{\mathrm{d}\mu_{\phi}}{\mathrm{d}\nu}(\xi)$ $=\underset{(\mathbf{x}_{u})_{u\in L}}{\mathbb{E}}\left[\ {\prod_{\begin{subarray}{c}u\in L\end{subarray}}\prod_{\begin{subarray}{c}v\in R\end{subarray}}\frac{\sigma_{\phi}(\langle\mathbf{x}_{u},\mathbf{x}_{v}\rangle)^{\xi_{u,v}}\left(1-\sigma_{\phi}(\langle\mathbf{x}_{u},\mathbf{x}_{v}\rangle)\right)^{1-\xi_{u,v}}}{p^{\xi_{u,v}}(1-p)^{1-\xi_{u,v}}}}\right]\text{ with }\sigma_{\phi}(\langle\mathbf{x}_{u},\mathbf{x}_{v}\rangle)\coloneqq\begin{cases}p&\text{if }\mathbf{M}_{u,v}=1\\ \sigma(\langle\mathbf{x}_{u},\mathbf{x}_{v}\rangle)&\text{otherwise}.\end{cases}$
To plug this back into our expression for $\chi^{2}(\mathcal{L}(\mu\mid S_{\rho}),\nu)$, we express our latent randomness as $\phi^{(1)}\coloneqq(\mathbf{X}_{\text{R}}^{(1)},\mathbf{M}^{(1)})$ and $\phi^{(2)}\coloneqq(\mathbf{X}_{\text{R}}^{(2)},\mathbf{M}^{(2)})$, and—in order to simplify notation—we abbreviate $\sigma^{(1)}_{1,v}=\sigma_{\phi^{(1)}}(\langle\mathbf{x}_{u}^{(1)},\mathbf{x}_{v}^{(1)}\rangle)$ and $\sigma^{(2)}_{1,v}=\\ \sigma_{\phi^{(2)}}(\langle\mathbf{x}_{u}^{(2)},\mathbf{x}_{v}^{(2)}\rangle)$. Then, the inner expectation evaluates as
$\underset{\xi\sim\nu}{\mathbb{E}}\left[\frac{\mathrm{d}\mu_{\phi^{(1)}}}{\mathrm{d}\nu}(\xi)\frac{\mathrm{d}\mu_{\phi^{(2)}}}{\mathrm{d}\nu}(\xi)\right]=\underset{(\mathbf{x}_{u}^{(1)},\mathbf{x}_{u}^{(2)})_{u\in L}}{\mathbb{E}}\left[\prod_{u\in L}\prod_{v\in R}\ \underset{\xi_{u,v}\sim\mathsf{Bern}(p)}{\mathbb{E}}\left[\left(\frac{1}{p^{2}}\sigma^{(1)}_{1,v}\sigma^{(2)}_{1,v}\right)^{\xi_{u,v}}\left(\frac{1}{(1-p)^{2}}\left(1-\sigma^{(1)}_{1,v}\right)\left(1-\sigma^{(2)}_{1,v}\right)\right)^{1-\xi_{u,v}}\right]\right]$
$\hskip 42.67912pt=\underset{\mathbf{x}_{1}^{(1)},\mathbf{x}_{1}^{(2)}}{\mathbb{E}}\left[\prod_{v\in R}\frac{1}{p}\sigma^{(1)}_{1,v}\sigma^{(2)}_{1,v}+\frac{1}{1-p}\left(1-\sigma^{(1)}_{1,v}\right)\left(1-\sigma^{(2)}_{1,v}\right)\right]^{m}$
@@
Expanding the above product, we get
$\underset{\xi\sim\nu}{\mathbb{E}}\left[\frac{\mathrm{d}\mu_{\phi^{(1)}}}{\mathrm{d}\nu}(\xi)\frac{\mathrm{d}\mu_{\phi^{(2)}}}{\mathrm{d}\nu}(\xi)\right]=\left(1+\sum_{\emptyset\neq\alpha\subseteq R}\frac{1}{(p(1-p))^{|\alpha|}}\prod_{u\in\alpha}\mathbb{E}_{\mathbf{x}_{u}}\left[(\sigma^{(1)}_{1,v}-p)\right]\mathbb{E}_{\mathbf{x}_{u}}\left[(\sigma^{(2)}_{1,v}-p)\right]\right)^{m}$
where we used that all the $(\mathbf{x}_{u}^{(1)},\mathbf{x}_{u}^{(2)})_{u\in L}$ are i.i.d. such that the product over $u\in L$ can be pulled out of the expectation and results in raising the entire expression to the power of $m=|L|$. Now, we can take the expectation over $\phi^{(1)},\phi^{(2)}$ and due to the independence of the entries in $\mathbf{M}^{(1)},\mathbf{M}^{(2)}$, we can take the expectation over $\mathbf{M}^{(1)},\mathbf{M}^{(2)}$ into the products above in order to obtain
@@
$\hskip 56.9055pt=\underset{\mathbf{X}_{\text{R}}^{(1)},\mathbf{X}_{\text{R}}^{(2)}\sim S_{\rho}}{\mathbb{E}}\left[\left(1+\sum_{\emptyset\neq\alpha\subseteq R}\frac{1}{(p(1-p))^{|\alpha|}}\prod_{u\in\alpha}\mathbb{E}_{\mathbf{x}_{u},\mathbf{M}^{(1)}_{u,v}}\left[(\sigma^{(1)}_{1,v}-p)\right]\mathbb{E}_{\mathbf{x}_{u},\mathbf{M}^{(2)}_{u,v}}\left[(\sigma^{(2)}_{1,v}-p)\right]\right)^{m}\right]$
@@
and where the inner expectations are over a single $\mathbf{x}\sim\mathcal{N}(0,\mathbf{I}_{d})$ and to be understood as conditional on $\mathbf{X}_{\text{R}}^{(1)},\mathbf{X}_{\text{R}}^{(2)}$777we omitted an explicit conditioning here and in the following in order to simplify notation.. Note that $\mathbb{E}\left[\textsc{Sw}^{(1)}(K_{1,\alpha})\right]$ is the expected signed weight of a star with set of leaves $\alpha$ conditional on the latent vectors of the leaves but over the randomness of the latent vector of the center and the mask $\mathbf{M}^{(1)}$.
@@ 5.2. The case of $p=1/2$
$=\underset{\mathbf{X}_{\text{R}}^{(1)},\mathbf{X}_{\text{R}}^{(2)}\sim S_{\rho}}{\mathbb{E}}\left[\left(1+\sum_{\emptyset\neq\alpha\subseteq R}\frac{1}{(p(1-p))^{|\alpha|}}\prod_{u\in\alpha}\mathbb{E}_{\mathbf{M}_{u,v}}\left[\mathbb{E}_{\mathbf{x}_{u}}[(\sigma^{(1)}_{1,v}-p)]\mathbb{E}_{\mathbf{x}_{u}}[(\sigma^{(2)}_{1,v}-p)]\right]\right)^{m}\right]$

### 15 Mar 2026 v4

@@ 5. Information-Theoretic Hardness
$\frac{\mathrm{d}\mu_{\phi}}{\mathrm{d}\nu}(\xi)$ $=\underset{(\mathbf{x}_{u})_{u\in L}}{\mathbb{E}}\left[\ {\prod_{\begin{subarray}{c}u\in L\end{subarray}}\prod_{\begin{subarray}{c}v\in R\end{subarray}}\frac{\sigma_{\phi}(\langle\mathbf{x}_{u},\mathbf{x}_{v}\rangle)^{\xi_{u,v}}\left(1-\sigma_{\phi}(\langle\mathbf{x}_{u},\mathbf{x}_{v}\rangle)\right)^{1-\xi_{u,v}}}{p^{\xi_{u,v}}(1-p)^{1-\xi_{u,v}}}}\right]\text{ with }\sigma_{\phi}(\langle\mathbf{x}_{u},\mathbf{x}_{v}\rangle)\coloneqq\begin{cases}p&\text{if }\mathbf{M}_{u,v}=0\\ \sigma(\langle\mathbf{x}_{u},\mathbf{x}_{v}\rangle)&\text{otherwise}.\end{cases}$
To plug this back into our expression for $\chi^{2}(\mathcal{L}(\mu\mid S_{\rho}),\nu)$, we express our latent randomness as $\phi^{(1)}\coloneqq(\mathbf{X}_{\text{R}}^{(1)},\mathbf{M}^{(1)})$ and $\phi^{(2)}\coloneqq(\mathbf{X}_{\text{R}}^{(2)},\mathbf{M}^{(2)})$, and—in order to simplify notation—we abbreviate $\sigma^{(1)}_{u,v}=\sigma_{\phi^{(1)}}(\langle\mathbf{x}_{u}^{(1)},\mathbf{x}_{v}^{(1)}\rangle)$ and $\sigma^{(2)}_{u,v}=\\ \sigma_{\phi^{(2)}}(\langle\mathbf{x}_{u}^{(2)},\mathbf{x}_{v}^{(2)}\rangle)$. Then, the inner expectation evaluates as
$\underset{\xi\sim\nu}{\mathbb{E}}\left[\frac{\mathrm{d}\mu_{\phi^{(1)}}}{\mathrm{d}\nu}(\xi)\frac{\mathrm{d}\mu_{\phi^{(2)}}}{\mathrm{d}\nu}(\xi)\right]=\underset{(\mathbf{x}_{u}^{(1)},\mathbf{x}_{u}^{(2)})_{u\in L}}{\mathbb{E}}\left[\prod_{u\in L}\prod_{v\in R}\ \underset{\xi_{u,v}\sim\mathsf{Bern}(p)}{\mathbb{E}}\left[\left(\frac{1}{p^{2}}\sigma^{(1)}_{u,v}\sigma^{(2)}_{u,v}\right)^{\xi_{u,v}}\left(\frac{1}{(1-p)^{2}}(1-\sigma^{(1)}_{u,v})(1-\sigma^{(2)}_{u,v})\right)^{1-\xi_{u,v}}\right]\right]$
$\hskip 42.67912pt=\underset{\mathbf{x}_{1}^{(1)},\mathbf{x}_{1}^{(2)}}{\mathbb{E}}\left[\prod_{v\in R}\frac{1}{p}\sigma^{(1)}_{1,v}\sigma^{(2)}_{1,v}+\frac{1}{1-p}(1-\sigma^{(1)}_{1,v})(1-\sigma^{(2)}_{1,v})\right]^{m}$
@@
where we used that all the $(\mathbf{x}_{u}^{(1)},\mathbf{x}_{u}^{(2)})_{u\in L}$ are i.i.d. such that the product over $u\in L$ can be pulled out of the expectation and results in raising the entire expression to the power of $m=|L|$. Expanding the above product, we get
$\underset{\xi\sim\nu}{\mathbb{E}}\left[\frac{\mathrm{d}\mu_{\phi^{(1)}}}{\mathrm{d}\nu}(\xi)\frac{\mathrm{d}\mu_{\phi^{(2)}}}{\mathrm{d}\nu}(\xi)\right]=\left(1+\sum_{\emptyset\neq\alpha\subseteq R}\frac{1}{(p(1-p))^{|\alpha|}}\mathbb{E}_{\mathbf{x}_{1}^{(1)}}\left[\prod_{v\in\alpha}(\sigma^{(1)}_{1,v}-p)\right]\mathbb{E}_{\mathbf{x}_{1}^{(2)}}\left[\prod_{v\in\alpha}(\sigma^{(2)}_{1,v}-p)\right]\right)^{m}$
Now, we can take the expectation over $\phi^{(1)},\phi^{(2)}$ and due to the independence of the entries in $\mathbf{M}^{(1)},\mathbf{M}^{(2)}$, we can take the expectation over $\mathbf{M}^{(1)},\mathbf{M}^{(2)}$ into the products above in order to obtain
@@
$\hskip 56.9055pt=\underset{\mathbf{X}_{\text{R}}^{(1)},\mathbf{X}_{\text{R}}^{(2)}\sim S_{\rho}}{\mathbb{E}}\left[\left(1+\sum_{\emptyset\neq\alpha\subseteq R}\frac{1}{(p(1-p))^{|\alpha|}}\mathbb{E}_{\mathbf{x}_{1}^{(1)}}\left[\prod_{u\in\alpha}\mathbb{E}_{\mathbf{M}^{(1)}_{1,v}}\left[(\sigma^{(1)}_{1,v}-p)\right]\right]\mathbb{E}_{\mathbf{x}_{1}^{(2)}}\left[\prod_{u\in\alpha}\mathbb{E}_{\mathbf{M}^{(2)}_{u,v}}\left[(\sigma^{(2)}_{1,v}-p)\right]\right]\right)^{m}\right]$
@@
and where we used
$\mathbb{E}_{\mathbf{M}^{(1)}_{1,v}}[(\sigma^{(1)}_{1,v}-p)]=q\sigma(\langle\mathbf{x}_{1}^{(1)},\mathbf{x}_{u}^{(1)}\rangle),\text{ and }\mathbb{E}_{\mathbf{M}^{(2)}_{1,v}}[(\sigma^{(2)}_{1,v}-p)]=q\sigma(\langle\mathbf{x}_{1}^{(2)},\mathbf{x}_{u}^{(2)}\rangle)$
. In the last line of (5), the inner expectations are over a single $\mathbf{x}\sim\mathcal{N}(0,\mathbf{I}_{d})$ and to be understood as conditional on $\mathbf{X}_{\text{R}}^{(1)},\mathbf{X}_{\text{R}}^{(2)}$777we omitted an explicit conditioning here and in the following in order to simplify notation.. Note that $\mathbb{E}\left[\textsc{Sw}^{(1)}(K_{1,\alpha})\right]$ is the expected signed weight of a star with set of leaves $\alpha$ conditional on the latent vectors of the leaves but over the randomness of the latent vector of the center and the mask $\mathbf{M}^{(1)}$.
@@ 5.2. The case of $p=1/2$
$=\underset{\mathbf{X}_{\text{R}}^{(1)},\mathbf{X}_{\text{R}}^{(2)}\sim S_{\rho}}{\mathbb{E}}\left[\left(1+\sum_{\emptyset\neq\alpha\subseteq R}\frac{1}{(p(1-p))^{|\alpha|}}\prod_{u\in\alpha}\mathbb{E}_{\mathbf{M}_{u,v}}\left[\mathbb{E}_{\mathbf{x}_{u}}[(\sigma^{(1)}_{u,v}-p)]\mathbb{E}_{\mathbf{x}_{u}}[(\sigma^{(2)}_{u,v}-p)]\right]\right)^{m}\right]$